# Shallow brambles

### v2

@@ Abstract.
A graph class $\mathcal{C}$ has polynomial expansion if there is a polynomial function $f$ such that for every graph $G\in\mathcal{C}$, each of the depth-$r$ minors of $G$ has average degree at most $f(r)$. In this note, we study bounded-radius variants of some classical graph parameters such as bramble number, linkedness and well-linkedness, and we show that they are pairwise polynomially related. Furthermore, in a monotone graph class with polynomial expansion they are all uniformly bounded by a polynomial in $r$.
@@ 1. Introduction
In other words, $\nabla_{r}(G)$ is half of the average degree of $H$, maximized over all depth-$r$ minors $H$ of $G$. A graph class $\mathcal{C}$ has bounded expansion if there is a function $f$ such that $\nabla_{r}(G)\leqslant f(r)$ for every $G\in\mathcal{C}$ and every integer $r\geqslant 0$. Furthermore, $\mathcal{C}$ has polynomial expansion if $f$ can be taken to be a polynomial.
@@
Moreover, $\mathrm{scol}_{1}(G)$ is equal to the coloring number of $G$ (one plus the degeneracy), whereas on the other side of the spectrum $\mathrm{scol}_{\infty}(G)$ equals the treewidth plus one. Zhu [19] proved that the following is equivalent for a class of graphs $\mathcal{C}$:
  1. (1)
@@
Let us now turn our attention to classes with polynomial expansion. Denote by $\omega_{r}(G)$ the size of a largest clique in a depth-$r$ minor of $G$. Then the following is equivalent for a class of graphs $\mathcal{C}$:
@@
Moreover, if $\mathcal{C}$ is monotone [^1] A class of graphs $\mathcal{C}$ is monotone if any subgraph of a graph from $\mathcal{C}$ is also in $\mathcal{G}$. then the following is also equivalent:
@@
The equivalence between (1) and (3) is due to Dvořák and Norin [3], while that of (1) and (2) has been shown in [11, Chapter 6]. Having balanced separators of size strongly sublinear in the number of vertices is a very useful property in structural graph theory and algorithm design, which motivates the study of graph classes with polynomial expansion. However, these classes are far from being well understood. Indeed, not much else is known besides the equivalences stated above.
@@
In this note, we introduce four new graph parameters that are bounded radius refinements of classical graph parameters: the $r$-bramble number, $r$-tangle number, $r$-linkedness and $r$-well-linkedness of a graph. Our main result is that, for some absolute constants $c_{1},c_{2}>0$, these new parameters are all lower bounded by polynomials in $r$ and $\omega_{c_{1}\cdot r}(G)$, and likewise they are all upper bounded by polynomials in $r$ and $\omega_{c_{2}\cdot r}(G)$; see Theorem 4. In particular, these parameters are all polynomially bounded in $r$ for graphs belonging to a fixed class with polynomial expansion.
It is our hope that these parameters might help in answering Question 1 positively. In particular, to do so it would be enough to show that, for graphs belonging to a fixed class with polynomial expansion, strong $r$-coloring numbers can be bounded from above by a polynomial function of $r$ and the $r$-bramble number:
@@
Let us emphasize the following subtlety in Question 2: The assumption that $G$ belongs to a fixed class with polynomial expansion is essential, because if we consider the set of all graphs, the $r$-strong coloring numbers cannot be bounded from above by any function of $r$ (polynomial or not) and the $r$-bramble number. This remains true even if we replace the $r$-bramble number by the $g(r)$-bramble number for any function $g(r)$, see 5 in Section 4.
@@
It was pointed out to us by Michał Pilipczuk that our definition of a shallow bramble is related to the cop-width of a graph, a parameter recently introduced by Toruńczyk [18]. Note that very much like brambles (or heavens) in the helicopter version of the cops and robber game of Seymour and Thomas [16], the existence of a radius-$r$ bramble of order $k$ provides a winning strategy for the robber against $k-1$ cops in the version of the game by Toruńczyk when the robber has speed $12r+$ 1.
@@ 2. Definitions and notation
A minor-model of a graph $H$ in a graph $G$ is a collection =$\mathcal{M}=\{M_{v}:v\in V(H)\}$ of disjoint subsets of $V(G)$, one for each vertex $v$ of $H$, such that $G[M_{v}]$ is connected for every $v\in V(H)$, and there is an edge in $G$ between $M_{v}$ and $M_{w}$ for every edge $vw\in E(H)$. If moreover $G[M_{v}]$ has radius at most $r$ for every $v\in V(H)$, we say that $\mathcal{M}$ is a depth-$r$ minor-model of $H$ in $G$.
Two vertex subsets of a graph $G$ are said to touch if they have a vertex in common or are joined by an edge of $G$. A bramble $\mathcal{B}$ in $G$ is a set of pairwise touching connected  [^3] Here we say that a subset $B\subseteq V(G)$ is connected if the induced subgraph $G[B]$ is. subsets of $V(G)$, called elements of $\mathcal{B}$. Note that a bramble can be viewed as a generalization of a minor-model of a complete graph, with the difference that its elements are allowed to have vertices in common. A hitting set of $\mathcal{B}$ is a set of vertices $X$ such $B\cap X\neq\emptyset$ for each $B\in\mathcal{B}$. The order of $\mathcal{B}$ is the minimum size of such a hitting set. Finally, the bramble number of $G$, denoted $\mathrm{bn}_{\infty}(G)$, is the maximum order of a bramble in $G$.
@@
  • For $r\in\mathbb{N}\cup\{\infty\}$, an $r$-bramble of a graph $G$ is a bramble such that each of its bramble elements induces a subgraph of $G$ of radius at most $r$. The maximum order of an $r$-bramble of $G$ is the $r$-bramble number of $G$, which we denote by $\mathrm{bn}_{r}(G)$.
  • For $t\geqslant 1$ and $r\in\mathbb{N}\cup\{\infty\}$, an $(r,t)$-bramble is an $r$-bramble $\mathcal{B}$ such that the intersection of any $t$ (not necessarily distinct) elements of $\mathcal{B}$ is non-empty. The maximum order of an $(r,t)$-bramble of a graph $G$ is the $(r,t)$-bramble number of $G$, denoted by $\mathrm{bn}_{r,t}(G)$.
The notion of $(r,t)$-bramble is of interest because for $t=3$ and $r=\infty$, it is closely related to a tangle. Tangles are usually defined in terms of separations, but as demonstrated in [13], the following yields an equivalent definition. For $r\in\mathbb{N}\cup\{\infty\}$, an $r$-tangle $\mathcal{T}$ of a graph $G$ is an $r$-bramble of $G$ such that for any three elements $T_{1},T_{2},T_{3}$ of $\mathcal{T}$, either
@@
The order of an $r$-tangle $\mathcal{T}$ is the order of $\mathcal{T}$ viewed as a bramble. The $r$-tangle number $\mathrm{tn}_{r}(G)$ of a graph $G$ is the maximum order of an $r$-tangle in $G$. The tangle number of $G$ is $\mathrm{tn}_{\infty}(G)$.
Observe that every $(r,3)$-bramble is an $r$-tangle, and every $r$-tangle is an $r$-bramble, so we have $\mathrm{bn}_{r,3}(G)\leqslant\mathrm{tn}_{r}(G)\leqslant\mathrm{bn}_{r}(G)$ for every graph $G$.
Let $r\in\mathbb{N}\cup\{\infty\}$ and let $k$ be an integer with $k\geqslant 1$. We say that a vertex set $S$ in a graph $G$ is $(r,k)$-linked if for any set $X$ of fewer than $k$ vertices, there is a connected subset in $G-X$ of radius at most $r$ that contains more than half of the vertices of $S$. The $r$-linkedness of $G$, denoted $\mathrm{link}_{r}(G)$, is the largest integer $k$ for which $G$ contains an $(r,k)$-linked set. The linkedness of $G$ is $\mathrm{link}_{\infty}(G)$.
We say that a set $S$ of vertices is $r$-well-linked if for every two subsets $A$ and $B$ of $S$ with $|A|=|B|$ and for every vertex set $Y$ of size less than $|A|$, there exists an $A-B$ path of length $\leqslant r$ in $G-Y$. The $r$-well-linkedness of $G$, denoted $\mathrm{well}_{r}(G)$, is the size of the largest $r$-well-linked set in $G$. The well-linkedness of $G$ is $\mathrm{well}_{\infty}(G)$.
We note that in the case $r=\infty$, by Menger’s theorem, the notion of an $r$-well-linked set could be equivalently defined by requiring the existence of $|A|$ vertex-disjoint $A-B$ paths (of length at most $r$) for every two equal-sized subsets $A$ and $B$ of $S$. However, for finite $r$ this equivalence is far from correct, due to examples in [10].
@@ 3. Main results
    $\mathrm{bn}_{r}(G)\leqslant(5r+1)\cdot(\omega_{5r+2}(G))^{2}$;
@@ 4. The proofs
We start with the following lemma mentioned in the introduction, showing that if we consider the set of all graphs, the strong $r$-coloring numbers cannot be bounded from above by any function of $r$ and the $g(r)$-bramble number, for any function $g(r)$.
@@
Let $\mathcal{B}$ be an $s$-bramble in $G$. Since any two bramble elements touch and each bramble element has radius at most $s$, it follows that the graph $G(\mathcal{B})$ induced by the union of the elements of $\mathcal{B}$ has diameter at most $14s+$ 1. Since this is less than half the girth of $G$, $G(\mathcal{B})$ must induce a tree. Trees have treewidth $1$, so every bramble (in particular every $s$-bramble) has a hitting set of size at most $2$. On the other hand, any edge $uv$ gives rise to the $s$-bramble :=$\mathcal{B}:=\{u,v\}$ that has order $2$, so $\mathrm{bn}_{s}(G)=2$. ∎
@@
The first inequality holds because every depth-$r$ minor-model of a complete graph $K_{n}$ is also an $r$-bramble of order $n$.
For the second inequality, let $\mathcal{B}$ be an $r$-bramble of $G$ and let $\pi$ be an ordering of $V(G)$. For each bramble element $B\in\mathcal{B}$, let $v_{B}$ be the minimum vertex of $B$ w.r.t. $\pi$, and let $v^{*}$ be the maximum vertex among all $v_{B}$ w.r.t. $\pi$. More formally, $v_{B}$ is the vertex such that $\pi(v_{B})=\min_{w\in B}\pi(w)$ and $v^{*}$ is such that $\pi(v^{*})=\max_{B\in\mathcal{B}}\min_{w\in B}\pi(w)$. Since $\mathcal{B}$ is an $r$-bramble, each bramble element of $\mathcal{B}$ has diameter at most $2r$. Thus for every $B\in\mathcal{B}$ there is a path $P(v^{*},B)$ from $v^{*}$ to $v_{B}$ of length at most $14r+$ 1. Traversing $P(v^{*},B)$ starting from $v^{*}$, let $x_{B}$ be the first vertex such that $\pi(x_{B})\leqslant\pi(v^{*})$. (Note that $x_{B}$ indeed exists since $\pi(v_{B})\leqslant\pi(v^{*})$.) Then $X:=\bigcup_{B\in\mathcal{B}}x_{B}$ is both a hitting set of $\mathcal{B}$ and a subset of $\mathrm{SReach}[4r+1,\pi,v^{*}]$. Since we derived this for arbitrary $\pi$ and $\mathcal{B}$, it follows that $\mathrm{bn}_{r}(G)\leqslant\mathrm{scol}_{4r+1}(G)$. ∎
@@
$\mathrm{bn}_{r}(G)\leqslant(5r+1)\cdot(\omega_{5r+2}(G))^{2}.$
@@
We prove that $\mathrm{bn}_{r}(G)\leqslant(5r+1)\cdot t^{2}$, where $t:=\omega_{5r+2}(G)$. Let $\mathcal{B}$ be an $r$-bramble of $G$. We call $\mathcal{M}$ a good model if
  • $\mathcal{M}$ is a depth-$(5r+2)$ minor-model of a complete graph, and
  • $|V(M)|\leqslant 1+5r(t-1)$ for every branch set $M\in\mathcal{M}$.
Note that there exists at least one good model, since a single vertex defines one. Given a good model $\mathcal{M}$, let $R(\mathcal{M})$ denote the union of all bramble elements of $\mathcal{B}$ that are disjoint from $V(\mathcal{M})$. Observe that, since every two bramble elements touch, the subgraph induced by $R(\mathcal{M})$ has radius at most $13r+$ 1.
From now on, we fix a good $\mathcal{M}$ such that $(|R(\mathcal{M})|,|V(\mathcal{M})|)$ is lexicographically minimized among all good models. Let $M_{1},M_{2},\ldots,M_{s}$ be the branch sets of $\mathcal{M}$. Note that by the definition of $t$, we have $s\leqslant t$.
The proof is done if $R(\mathcal{M})=\emptyset$, because then $V(\mathcal{M})$ is a hitting set of $\mathcal{B}$ of order at most $\sum_{1=1}^{s}|V(M_{i})|\leqslant t+5r(t-1)t\leqslant(5r+1)\cdot t^{2}$. Thus we may assume that $R(\mathcal{M})$ is non-empty, and from there we will derive a contradiction.
@@
for otherwise we would have $|R(\mathcal{M}-M_{i})|=|R(\mathcal{M})|$ and $|V(\mathcal{M}-M_{i})|<|V(\mathcal{M})|$, contradicting the minimality of $\mathcal{M}$.
Thanks to (1) we know that each branch set $M_{i}\in\mathcal{M}$ satisfies the following key property: There is a bramble element $B\in\mathcal{B}$ that intersects $M_{i}$ and avoids all other bramble elements. In $M_{i}\cup B$, take a shortest path $P_{i}$ from $M_{i}$ to the non-empty set $R(\mathcal{M})$. Let $x_{i}$ be the endpoint of $P_{i}$ in $R(\mathcal{M})$. Observe that $|P_{i}|\leqslant 2r+2$. Let $Q_{i}$ be $P_{i}$ minus its endpoint in $M_{i}$.
Now define a new branch set $M_{s+1}$ as follows: First take a rooted tree of radius at most $13r+$ 1 in $R(\mathcal{M})$ having $x_{1},\ldots,x_{s}$ as leaves, and then add the paths $Q_{1},\ldots,Q_{s}$. Then $\mathrm{rad}(M_{s+1})\leqslant 5r+2$, and $|V(M_{s+1})|\leqslant 1+5r(t-1)$. Moreover, by construction $M_{s+1}$ is connected to all previous branch sets. We conclude that $\mathcal{M}+M_{s+1}$ is a good model with $|R(\mathcal{M}+M_{s+1})|<|R(\mathcal{M})|$, contradicting the minimality of $\mathcal{M}$. ∎
@@
Let $S$ be an $(r,k)$-linked set, where $k:=\mathrm{link}_{r}(G)$. By definition of $(r,k)$-linkedness, for each set $X\subseteq V(G)$ of order $|X|<k$, there exists an $r$-ball $B_{X}$ in $G-X$ that contains more than half of the vertices of $S$. Let $\mathcal{B}:=(B_{X})_{X\subseteq V(G),|X|<k}$ be the set of these balls. The elements of $\mathcal{B}$ pairwise intersect since each of them contains more than half of $S$, and therefore $\mathcal{B}$ is an $r$-bramble of $G$. Furthermore, the order of $\mathcal{B}$ is at least $k$, since any hypothetical hitting set $X$ of order less than $k$ would not hit the bramble element $B_{X}$. Thus $\mathrm{bn}_{r}(G)\geqslant k$, as desired. ∎
@@
Let $S$ be a minimum order hitting set for an $r$-bramble $\mathcal{B}$. It suffices to show that $S$ is $(3r+1,\lceil|S|/2\rceil)-$linked. To that end, let $Y$ be a set of fewer than $\lceil|S|/2\rceil$ vertices.
Since $Y$, being smaller than $S$, is not a hitting set for $\mathcal{B}$, there exists a bramble element $B\in\mathcal{B}$ which is disjoint from $Y$. Because $B$ is an element of an $r$-bramble, we can choose a vertex $u$ in $B$ such that (in the graph induced by $B$ and thus also in $G-Y$) every vertex of $B$ is at distance at most $r$ from $u$. We let $U$ denote the $(3r+1)$-ball in $G-Y$ centered at $u$. It suffices to show that $U$ contains more than half of $S$, which we will now do.
@@
We prove that any minimum hitting set $S$ for an $r$-bramble $\mathcal{B}$ is $(4r+1)$-well-linked. Suppose not, then there exist two equal-sized subsets $A,B$ of $S$ and a set $Y$ of fewer than $|A|$ vertices, such that $G-Y$ does not contain any $A-B$ path of length $\leqslant 4r+1$. Consider the sub-bramble $\mathcal{B}_{Y}$ of $\mathcal{B}$ that contains all bramble elements of $\mathcal{B}$ that are disjoint from $Y$. If both $A$ and $B$ intersect some (not necessarily the same) bramble element of $\mathcal{B}_{Y}$ then these bramble elements touch, so then there must be a path of length at most $14r+$ 1 between $A$ and $B$ in $G-Y$, a contradiction. So without loss of generality $\bigcup_{C\in\mathcal{B}_{Y}}C\cap A$ is empty. But then $(S-A)+Y$ is a hitting set for $\mathcal{B}$, contradicting the minimality of $S$. ∎

### v3

@@ Abstract.
A graph class $\mathcal{C}$ has polynomial expansion if there is a polynomial function $f$ such that for every graph $G\in\mathcal{C}$, each of the depth-$r$ minors of $G$ has average degree at most $f(r)$. In this note, we study bounded-radius variants of some classical graph parameters such as bramble number, linkedness and well-linkedness, and we show that they are pairwise polynomially related. Furthermore, in a graph class with polynomial expansion they are all uniformly bounded by a polynomial in $r$.
@@ 1. Introduction
In other words, $\nabla_{r}(G)$ is the maximum edge density of a depth-$r$ minor of $G$. A graph class $\mathcal{C}$ has bounded expansion if there is a function $f$ such that $\nabla_{r}(G)\leqslant f(r)$ for every $G\in\mathcal{C}$ and every integer $r\geqslant 0$. Furthermore, $\mathcal{C}$ has polynomial expansion if $f$ can be taken to be a polynomial.
@@
Moreover, $\mathrm{scol}_{1}(G)$ is equal to the coloring number of $G$ (one plus the degeneracy), whereas on the other side of the spectrum $\mathrm{scol}_{\infty}(G)$ equals the treewidth plus one. Zhu [19] proved that the following statements are equivalent for a class of graphs $\mathcal{C}$:
  1. (i)
@@
Let us now turn our attention to classes with polynomial expansion. Denote by $\omega_{r}(G)$ the size of a largest clique in a depth-$r$ minor of $G$. Then the following statements are equivalent for a class of graphs $\mathcal{C}$ [11]:
@@
Moreover, if $\mathcal{C}$ is monotone [^1] A class of graphs $\mathcal{C}$ is monotone if any subgraph of a graph from $\mathcal{C}$ is also in $\mathcal{C}$. then the following is also equivalent to (1) and (2) (see [3]):
@@
Having balanced separators of size strongly sublinear in the number of vertices is a very useful property in structural graph theory and algorithm design, which motivates the study of graph classes with polynomial expansion. However, these classes are far from being well understood. Indeed, not much else is known besides the equivalences stated above.
@@
In this note, we introduce four new graph parameters that are bounded radius refinements of classical graph parameters: the depth-$r$ bramble number, depth-$r$ tangle number, depth-$r$ linkedness and depth-$r$ well-linkedness of a graph. All these parameters quantify how well-connected a subset of vertices of bounded diameter can be in a given graph. Our main result is that, for some absolute constants $c_{1},c_{2}>0$, these new parameters are all lower bounded by polynomials in $r$ and $\omega_{c_{1}\cdot r}(G)$, and likewise they are all upper bounded by polynomials in $r$ and $\omega_{c_{2}\cdot r}(G)$; see Theorem 4. In particular, these parameters are all polynomially bounded in $r$ for graphs belonging to a fixed class with polynomial expansion.
It is our hope that these parameters might help in answering Question 1 positively. In particular, to do so it would be enough to show that, for graphs belonging to a fixed class with polynomial expansion, strong $r$-coloring numbers can be bounded from above by a polynomial function of $r$ and the depth-$r$ bramble number:
@@
Let us emphasize the following subtlety in Question 2: The assumption that $G$ belongs to a fixed class with polynomial expansion is essential, because if we consider the set of all graphs, the $r$-strong coloring numbers cannot be bounded from above by any function of $r$ (polynomial or not) and the depth-$r$ bramble number. This remains true even if we replace the depth-$r$ bramble number by the depth-$g(r)$ bramble number for any function $g(r)$, see 5 in Section 4.
@@
It was pointed out to us by Michał Pilipczuk that our definition of a shallow bramble is related to the cop-width of a graph, a parameter recently introduced by Toruńczyk [18]. Note that very much like brambles (or heavens) in the helicopter version of the cops and robber game of Seymour and Thomas [16], the existence of a depth-$r$ bramble of order $k$ provides a winning strategy for the robber against $k-1$ cops in the version of the game by Toruńczyk when the robber has speed $12r+$ 1.
@@ 2. Definitions and notation
We say that a subset $B\subseteq V(G)$ is connected if the induced subgraph $G[B]$ is. A minor-model of a graph $H$ in a graph $G$ is a collection =$\mathcal{M}=\{M_{v}:v\in V(H)\}$ of pairwise disjoint connected subsets of $V(G)$, one for each vertex $v$ of $H$, such that there is an edge in $G$ between $M_{v}$ and $M_{w}$ for every edge $vw\in E(H)$. If moreover $G[M_{v}]$ has radius at most $r$ for every $v\in V(H)$, we say that $\mathcal{M}$ is a depth-$r$ minor-model of $H$ in $G$.
Two vertex subsets of a graph $G$ are said to touch if they have a vertex in common or are joined by an edge of $G$. A bramble $\mathcal{B}$ in $G$ is a set of pairwise touching connected subsets of $V(G)$, called elements of $\mathcal{B}$. Note that a bramble can be viewed as a generalization of a minor-model of a complete graph, with the difference that its elements are allowed to have vertices in common. A hitting set of $\mathcal{B}$ is a set of vertices $X$ such that $B\cap X\neq\emptyset$ for each $B\in\mathcal{B}$. The order of $\mathcal{B}$ is the minimum size of such a hitting set. Finally, the bramble number of $G$, denoted $\mathrm{bn}_{\infty}(G)$, is the maximum order of a bramble in $G$. A classical result of Seymour and Thomas [16] states that $\mathrm{bn}_{\infty}(G)$ is equal to the treewidth of $G$ plus one.
@@
  • For $r\in\mathbb{N}\cup\{\infty\}$, a depth-$r$ bramble of a graph $G$ is a bramble such that each of its bramble elements induces a subgraph of $G$ of radius at most $r$. The maximum order of a depth-$r$ bramble of $G$ is the depth-$r$ bramble number of $G$, which we denote by $\mathrm{bn}_{r}(G)$.
  • For $t\geqslant 1$ and $r\in\mathbb{N}\cup\{\infty\}$, a depth-$r$ $t$-bramble is a depth-$r$ bramble $\mathcal{B}$ such that the intersection of any $t$ (not necessarily distinct) elements of $\mathcal{B}$ is non-empty. The maximum order of a depth-$r$ $t$-bramble of a graph $G$ is the depth-$r$ $t$-bramble number of $G$, denoted by $\mathrm{bn}_{r,t}(G)$.
The notion of depth-$r$ $t$-bramble is of interest because for $t=3$ and $r=\infty$, it is closely related to a tangle. Tangles are usually defined in terms of separations, but as demonstrated in [13], the following yields an equivalent definition. For $r\in\mathbb{N}\cup\{\infty\}$, a depth-$r$ tangle $\mathcal{T}$ of a graph $G$ is a depth-$r$ bramble of $G$ such that for any three (not necessarily distinct) elements $T_{1},T_{2},T_{3}$ of $\mathcal{T}$, either
@@
The order of a depth-$r$ tangle $\mathcal{T}$ is the order of $\mathcal{T}$ viewed as a bramble. The depth-$r$ tangle number $\mathrm{tn}_{r}(G)$ of a graph $G$ is the maximum order of a depth-$r$ tangle in $G$. The tangle number of $G$ is $\mathrm{tn}_{\infty}(G)$.
Observe that every depth-$r$ $3$-bramble is a depth-$r$ tangle, and every depth-$r$ tangle is a depth-$r$ bramble, so we have $\mathrm{bn}_{r,3}(G)\leqslant\mathrm{tn}_{r}(G)\leqslant\mathrm{bn}_{r}(G)$ for every graph $G$.
Let $r\in\mathbb{N}\cup\{\infty\}$ and let $k$ be an integer with $k\geqslant 1$. We say that a vertex set $S$ in a graph $G$ is depth-$r$ $k$-linked if for any set $X$ of fewer than $k$ vertices, there is a connected subset in $G-X$ of radius at most $r$ that contains more than half of the vertices of $S$. The depth-$r$ linkedness of $G$, denoted $\mathrm{link}_{r}(G)$, is the largest integer $k$ for which $G$ contains a depth-$r$ $k$-linked set. The linkedness of $G$ is $\mathrm{link}_{\infty}(G)$.
We say that a set $S$ of vertices is depth-$r$ well-linked if for every two subsets $A$ and $B$ of $S$ with $|A|=|B|$ and for every vertex set $Y$ of size less than $|A|$, there exists an $A-B$ path of length $\leqslant r$ in $G-Y$. The depth-$r$ well-linkedness of $G$, denoted $\mathrm{well}_{r}(G)$, is the size of the largest depth-$r$ well-linked set in $G$. The well-linkedness of $G$ is $\mathrm{well}_{\infty}(G)$.
We note that in the case $r=\infty$, by Menger’s theorem, the notion of a depth-$r$ well-linked set could be equivalently defined by requiring the existence of $|A|$ vertex-disjoint $A-B$ paths (of length at most $r$) for every two equal-sized subsets $A$ and $B$ of $S$. However, for finite $r$ this equivalence is far from correct, due to examples in [10].
@@ 3. Main results
    $\mathrm{bn}_{r}(G)\leqslant(5r+1)\cdot(\omega_{5r+1}(G))^{2}$;
@@ 4. The proofs
We start with the following lemma mentioned in the introduction, showing that if we consider the set of all graphs, the strong $r$-coloring numbers cannot be bounded from above by any function of $r$ and the depth-$g(r)$ bramble number, for any function $g(r)$.
@@
Let $\mathcal{B}$ be a depth-$s$ bramble in $G$. Since any two bramble elements touch and each bramble element has radius at most $s$, it follows that the graph $G(\mathcal{B})$ induced by the union of the elements of $\mathcal{B}$ has diameter at most $14s+$ 1. Since this is less than half the girth of $G$, $G(\mathcal{B})$ must induce a tree. Trees have treewidth $1$, so every bramble (in particular every depth-$s$ bramble) has a hitting set of size at most $2$ [16]. On the other hand, any edge $uv$ gives rise to the depth-$s$ bramble :=$\mathcal{B}:=\{u,v\}$ that has order $2$, so $\mathrm{bn}_{s}(G)=2$. ∎
@@
The first inequality holds because every depth-$r$ minor-model of a complete graph $K_{n}$ is also a depth-$r$ bramble of order $n$.
For the second inequality, let $\mathcal{B}$ be a depth-$r$ bramble of $G$ and let $\pi$ be an ordering of $V(G)$. For each bramble element $B\in\mathcal{B}$, let $v_{B}$ be the minimum vertex of $B$ w.r.t. $\pi$, and let $v^{*}$ be the maximum vertex among all $v_{B}$ w.r.t. $\pi$. More formally, $v_{B}$ is the vertex such that $\pi(v_{B})=\min_{w\in B}\pi(w)$ and $v^{*}$ is such that $\pi(v^{*})=\max_{B\in\mathcal{B}}\min_{w\in B}\pi(w)$. Since $\mathcal{B}$ is a depth-$r$ bramble, each bramble element of $\mathcal{B}$ has diameter at most $2r$. Thus for every $B\in\mathcal{B}$ there is a path $P(v^{*},B)$ from $v^{*}$ to $v_{B}$ of length at most $14r+$ 1. Traversing $P(v^{*},B)$ starting from $v^{*}$, let $x_{B}$ be the first vertex such that $\pi(x_{B})\leqslant\pi(v^{*})$. (Note that $x_{B}$ indeed exists since $\pi(v_{B})\leqslant\pi(v^{*})$.) Then $X:=\bigcup_{B\in\mathcal{B}}x_{B}$ is both a hitting set of $\mathcal{B}$ and a subset of $\mathrm{SReach}[4r+1,\pi,v^{*}]$. Since we derived this for arbitrary $\pi$ and $\mathcal{B}$, it follows that $\mathrm{bn}_{r}(G)\leqslant\mathrm{scol}_{4r+1}(G)$. ∎
@@
$\mathrm{bn}_{r}(G)\leqslant(5r+1)\cdot(\omega_{5r+1}(G))^{2}.$
@@
We prove that $\mathrm{bn}_{r}(G)\leqslant(5r+1)\cdot t^{2}$, where $t:=\omega_{5r+1}(G)$. Let $\mathcal{B}$ be a depth-$r$ bramble of $G$. We call $\mathcal{M}$ a good model if
  • $\mathcal{M}$ is a depth-$(5r+1)$ minor-model of a complete graph in $G$, and
  • $|V(M)|\leqslant 1+5(r+1)(t-1)$ for every branch set $M\in\mathcal{M}$.
Note that there exists at least one good model, since a single vertex defines one. Given a minor-model $\mathcal{M}$, let $V(\mathcal{M})$ denote the union of the branch sets of $\mathcal{M}$, and let $R(\mathcal{M})$ denote the union of all bramble elements of $\mathcal{B}$ that are disjoint from $V(\mathcal{M})$. Observe that, since every two bramble elements touch, the subgraph induced by $R(\mathcal{M})$ has radius at most $13r+$ 1.
From now on, we fix a good model $\mathcal{M}$ in $G$ such that $(|R(\mathcal{M})|,|V(\mathcal{M})|)$ is lexicographically minimized among all good models. Let $M_{1},M_{2},\ldots,M_{s}$ be the branch sets of $\mathcal{M}$. Note that by the definition of $t$, we have $s\leqslant t$.
The proof is done if $R(\mathcal{M})=\emptyset$, because then $V(\mathcal{M})$ is a hitting set of $\mathcal{B}$ of order at most $\sum_{1=1}^{s}|V(M_{i})|\leqslant t+(5r+1)(t-1)t\leqslant(5r+1)\cdot t^{2}$. Thus we may assume that $R(\mathcal{M})$ is non-empty, and from there we will derive a contradiction.
@@
for otherwise we would have $|R(\mathcal{M}-M_{i})|\leqslant|R(\mathcal{M})|$ and $|V(\mathcal{M}-M_{i})|<|V(\mathcal{M})|$, contradicting the minimality of $\mathcal{M}$.
Thanks to (1) we know that each branch set $M_{i}\in\mathcal{M}$ satisfies the following key property: There is a bramble element $B\in\mathcal{B}$ that intersects $M_{i}$ and avoids all other branch sets of $\mathcal{M}$. Take a shortest path $P_{i}$ from $M_{i}$ to the non-empty set $R(\mathcal{M})$ through $M_{i}\cup B$. Let $x_{i}$ be the endpoint of $P_{i}$ in $R(\mathcal{M})$. Observe that $P_{i}$ contains at most $12r+$ 1 edges, and thus at most $22r+$ 2 vertices. Let $Q_{i}$ be $P_{i}$ minus its endpoint in $M_{i}$ (so $Q_{i}$ contains at most $2r$ edges).
Recall that the subgraph induced by $R(\mathcal{M})$ has radius at most $13r+$ 1, so there exists a vertex $x^{*}\in R(\mathcal{M})$ and for each $1\leqslant i\leqslant s$, a path on at most $13r+$ 1 edges, included in $R(\mathcal{M})$ and connecting $x^{*}$ to $x_{i}$. Let $X\subseteq R(\mathcal{M})$ be the union of the vertex sets of these $s$ paths. Note that $X$ is connected and contains at most $1+(3r+1)s$ vertices.
Now define a new branch set $M_{s+1}$, consisting of $X$ together with the paths $Q_{1},\ldots,Q_{s}$. Then the radius of $M_{s+1}$ is at most $15r+$ 1, and $|V(M_{s+1})|\leqslant 1+(5r+1)s\leqslant 1+(5r+1)(t-1)$. Moreover, by construction $M_{s+1}$ is connected to all previous branch sets. We conclude that $\mathcal{M}+M_{s+1}$ is a good model with $|R(\mathcal{M}+M_{s+1})|<|R(\mathcal{M})|$, contradicting the minimality of $\mathcal{M}$. ∎
@@
Let $S$ be a depth-$r$ $k$-linked set, where $k:=\mathrm{link}_{r}(G)$. By definition of depth-$r$ $k$-linkedness, for each set $X\subseteq V(G)$ of order $|X|<k$, there exists an $r$-ball $B_{X}$ in $G-X$ that contains more than half of the vertices of $S$. Let $\mathcal{B}:=(B_{X})_{X\subseteq V(G),|X|<k}$ be the set of these balls. The elements of $\mathcal{B}$ pairwise intersect since each of them contains more than half of $S$, and therefore $\mathcal{B}$ is a depth-$r$ bramble of $G$. Furthermore, the order of $\mathcal{B}$ is at least $k$, since any hypothetical hitting set $X$ of order less than $k$ would not hit the bramble element $B_{X}$. Thus $\mathrm{bn}_{r}(G)\geqslant k$, as desired. ∎
@@
Let $S$ be a minimum order hitting set for a depth-$r$ bramble $\mathcal{B}$. It suffices to show that $S$ is depth-$(3r+1)$ $(\lceil|S|/2\rceil)-$linked. To that end, let $Y$ be a set of fewer than $\lceil|S|/2\rceil$ vertices.
Since $Y$, being smaller than $S$, is not a hitting set for $\mathcal{B}$, there exists a bramble element $B\in\mathcal{B}$ which is disjoint from $Y$. Because $B$ is an element of a depth-$r$ bramble, we can choose a vertex $u$ in $B$ such that (in the graph induced by $B$ and thus also in $G-Y$) every vertex of $B$ is at distance at most $r$ from $u$. We let $U$ denote the $(3r+1)$-ball in $G-Y$ centered at $u$. It suffices to show that $U$ contains more than half of $S$, which we will now do.
@@
We prove that any minimum hitting set $S$ for a depth-$r$ bramble $\mathcal{B}$ is depth-$(4r+1)$ well-linked. Suppose not, then there exist two equal-sized subsets $A,B$ of $S$ and a set $Y$ of fewer than $|A|$ vertices, such that $G-Y$ does not contain any $A-B$ path of length $\leqslant 4r+1$. Consider the sub-bramble $\mathcal{B}_{Y}$ of $\mathcal{B}$ that contains all bramble elements of $\mathcal{B}$ that are disjoint from $Y$. If both $A$ and $B$ intersect some (not necessarily the same) bramble element of $\mathcal{B}_{Y}$ then these bramble elements touch, so then there must be a path of length at most $14r+$ 1 between $A$ and $B$ in $G-Y$, a contradiction. So without loss of generality $\bigcup_{C\in\mathcal{B}_{Y}}C\cap A$ is empty. But then $(S-A)+Y$ is a hitting set for $\mathcal{B}$, contradicting the minimality of $S$. ∎